Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 28     Entry time: 21 Feb 2024, 23:00
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery: stable and flat over 250 pulses per train. 1.1 MHz with 400 uJ

Optical laser delivery: NA

Achievements/Observations:

  • Beam optimization and characterization
    - Short pulses tuned
    - HIREX setup and used to characterize the number of SASE SPIKES per pulse (~10)

Issues

  • 10 SASE spikes per pulse is more than desired
    - discussion with BKR and AccRC regarding possibilities for optimization (feedback in the morning)
    -HIREX data proc is not correct as dark processing fails

    - DOC informed and working on it
    - one issue was gothard II appeard to have timed out - does the number of triggers need increasing?
    -sample jetting setup

    - electrospray working but saw exactly 1 hit in run 32
    - tried to alighn with silver nano-cubes but had to switch to sucrose to increase hit rate to aid alignment
    - no hits with sucrose seen with scanning X +- 1mm

ELOG V3.1.4-7c3fd00